solve this system for x and y
(x+2/2)-(y-4/3)=4
(x+y/2)=(1/2)+(x-y/3)

Found 2 solutions by stanbon, richwmiller:
Answer by stanbon(75887)

solve this system for x and y
(x+2/2)-(y-4/3)=4
(x+y/2)=(1/2)+(x-y/3)
----------------------------------
Multiply thru the 1st by 6
Multiply thru the 2nd by 6
---------------------------------
3(x+2)-2(y-4) = 24
3(x+y) = 3 + 2(x-y)
--------------------------------
Rearrange:
3x - 2y = 10
x - 5y = 3
-----------------------
Solve for x by substitution or eliminination.
Then solve for y.
========================
Cheers,
Stan H.

Answer by richwmiller(17219)

((x+2)/2)-((y-4)/3)=4
((x+y)/2)=(1/2)+((x-y)/3)
multiply through by 6
3*(x+2)-2*(y-4)=24
3*(x+y)=3+2(x-y)
3x+6-2y+8=24
3x+3y=3+2x-2y
3x-2y=10
x=3-5y
3*(3-5y)-2y=10
9-15y-2y=10
-17y=1
x = 56/17, y = -1/17